OpenStudy (anonymous):

what is the simplified form of √605w^24?

OpenStudy (callisto):

Is the question \(\sqrt{605w^{24}}\)?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i suck at this. so i have no idea what to do.

OpenStudy (callisto):

605 = 5 x 11 x 11 = 5 x 11^2 \[\sqrt{605w^{24}} = \sqrt{5\times 11^2 w^{24}} = ...?\]

OpenStudy (anonymous):

are you asking me to solve or what?

OpenStudy (callisto):

Solve~

OpenStudy (anonymous):

ok um 5^4√11?

OpenStudy (anonymous):

just a guess based off of my answer choices.

OpenStudy (callisto):

\[\sqrt{5}\times \sqrt{11^2} \times \sqrt{w^{24}}= ...?\]

OpenStudy (anonymous):

i have no idea...

OpenStudy (callisto):

\[\sqrt{a^2} = a\]

OpenStudy (anonymous):

so 11^12√5